When r = 0, the best estimate of variable Y from a known variable X is…

a.            the standard deviation of X

b.            the standard deviation of Y

c.            Mean of X

d.            Mean of Y
Answer #1

Linear regression equation is - Y = b0 + b1X.

Where, b0 = \bar{y} - b_{1}\bar{x} & b1 = r\tfrac{s_{y}}{s_{x}}

Here, given that r = 0.

So, it implies that value of b1 also becomes 0.

Now, we have b0 = \bar{y} - b_{1}\bar{x} = \bar{y}-0(\bar{x}) = \bar{y}

Hence, linear regression equation becomes as -

Y = b0 + 0(X)= b0 = \bar{y} ........ ( since, b1 = 0)

Hence, best estimate of variable Y from known variable X is mean of y.

Option D is correct.